OpenStudy (adll23):

Math question The distance traveled by an object can be modeled by the equation d = ut + 0.5at2 where d = distance, u = initial velocity, t = time, and a = acceleration. Solve this formula for a. Show all steps in your work.

OpenStudy (amistre64):

its basic algebra .... subtract and divide

OpenStudy (adll23):

I'm not good at algebra so I came here for help

OpenStudy (amistre64):

not much to help with on this .. subtract, and then divide. where is the issue?

OpenStudy (adll23):

I just don't get what its asking lol

OpenStudy (amistre64):

d = ut + 0.5 a t2 ^ get this guy all by his little lonesome over here

OpenStudy (adll23):

k so I minus 0.5

OpenStudy (amistre64):

no, minus ut, since its just hanging about there.

OpenStudy (adll23):

k so how do I minus ut from d?

OpenStudy (amistre64):

d = ut + 0.5 a t2 -ut -ut ---------------- d-ut = .5 a t^2 now all tha tis left is sticking to the a by multiplication, so we divide it off

OpenStudy (amistre64):

just subtract ut from d ...

OpenStudy (adll23):

so then I divide what?

OpenStudy (amistre64):

whatever is not a

OpenStudy (amistre64):

.5 and t^2 are not a, so they can divide off

OpenStudy (adll23):

so I divide .5

OpenStudy (amistre64):

(d-ut) = (.5 a t^2) ---- -------- .5t^2 .5 t^2 since .5/.5 = 1 and t^2/t^2 = 1, a*1*1 = a all by itself

OpenStudy (amistre64):

as long as t is not 0, we should be fine

OpenStudy (adll23):

so the answer is a*1*1?

OpenStudy (amistre64):

no

OpenStudy (amistre64):

if a is all by itself on the right, then it equals all the is on the left ...

OpenStudy (adll23):

but you said a*1*1=a all by itself

OpenStudy (amistre64):

follow thru again ... d = ut + .5 a t^2 what should we do first?

OpenStudy (adll23):

minus ut from d

OpenStudy (adll23):

which would be ut-d=.5 a t^2

OpenStudy (amistre64):

and yes, i do see where that choice of wording was not a good choice on my part :) minus ut, from both sides. d-ut = .5 a t^2 notice how we are unwrapping the 'a' ... what next?

OpenStudy (adll23):

divide .5 and t^2

OpenStudy (amistre64):

good (d-ut)/(.5t^2) = a now 'a' is all alone, and it equals whatever is on the other side

OpenStudy (adll23):

so a equals (d-ut)/(.5t^d) or do we have to simplify it?

OpenStudy (amistre64):

im not grading it, so however farther you wish to take it is up to you.

OpenStudy (amistre64):

we got 'a' all by itself on one side ... so whatever else you do is no matter to me or to the math of 'a' ... it just makes the other side look differently

OpenStudy (adll23):

ok thank you!

OpenStudy (amistre64):

good luck :)